OpenStudy (anonymous):

simplify (x^2-16)/(x^2-3x-4)

OpenStudy (anonymous):

factor top and bottom

OpenStudy (anonymous):

\[\frac{x^2-16}{x^2-3x-4}=\frac{(x+4)(x-4)}{(x+1)(x-4)}\]

OpenStudy (anonymous):

so cancel to get \[\frac{x+4}{x+1}\]

OpenStudy (anonymous):

ty